# Discovering Concept Directions from Diffusion-based Counterfactuals via Latent Clustering

### v1

@@ 4.2 Generation of Counterfactuals
The same classifier-guided counterfactual generation approach introduced in the CDCT framework [15] is adopted, with a simplified setting: rather than capturing intermediate reconstructions across the trajectory, only the final counterfactual image is retained. Counterfactuals are generated using a Latent Diffusion Model based on the Stable Diffusion (SD) 2.1 architecture [^2] Model available at: https://huggingface.co/stabilityai/stable-diffusion-2-1., which is fine-tuned on the consolidated ISIC training dataset to align the generative manifold with the dermoscopic image domain. For each training sample, counterfactuals are synthesized for all target classes other than the class predicted by the classifier.
@@ 4.3 Extraction of Concept Directions
Semantic concept directions induced by classifier-guided counterfactuals are explored in two distinct latent spaces: (1) the pretrained encoder of the Stable Diffusion 2.1 model, referred to as the LDM encoder, and (2) a Variational Autoencoder trained on counterfactual trajectories, referred to as the CDCT encoder (see Step 2 of the CDCT framework [15] for further details). Experimental setup, results, and analysis based on the CDCT encoder are provided in the supplementary material.
@@
The unit vectors are collected for all training samples that are not predicted as a given target class. Spherical K-Means clustering [29] is then applied to group these vectors that reflect similar semantic transformations. The number of clusters is selected based on the silhouette coefficient to ensure cohesion and separation. Each resulting cluster represents a set of consistent latent transformations, and the average unit direction within a cluster is interpreted as a representative concept direction for the target class.
@@ 4.4 Interpretation and Evaluation of Concept Directions
To assess the interpretability and effectiveness of the extracted concept directions, each identified concept direction is applied to unseen test images in the latent space. Figure 2 illustrates concept directions identified by CDLC using the LDM encoder. The success rate reported in each sub-caption indicates the proportion of test samples for which traversal along the corresponding direction increases the predicted probability of the target class, relative to the reconstructed image.
For Melanoma, the extracted direction (Figure 2a) increases central pigmentation and introduces a purplish hue in the core lesion. In Basal Cell Carcinoma, detected concept direction (Figure 2b) produces a reddish-purple hue with blue-gray dots. For Benign Keratosis, a concept direction (Figure 2c) induces purplish blotches, irregular pigmentation, and blurred borders. In the case of Dermatofibroma, the direction (Figure 2d) introduces a violet-toned center with subtle border irregularities and increased asymmetric pigmentation. For Vascular Lesions, the direction (Figure 2e) enhances central coloration with purplish tones and peripheral pink diffusion. Finally, Squamous Cell Carcinoma directions (Figure 2f) reveal white central structures accompanied by irregular linear streaks suggestive of vascular features.
@@ 5 Discussion
The results demonstrate that CDLC effectively extracts global concept directions for skin lesion classification while significantly improving computational efficiency over CDCT. By avoiding iterative traversal of individual latent dimensions, CDLC reduces the overhead typically associated with concept discovery, especially in high-dimensional latent spaces. Additionally, it captures rich, multidimensional semantic directions beyond the axis-aligned perturbations of CDCT, which tend to detect only simple features. Beyond efficiency, CDLC reliably discovers class-specific concept directions by constructing latent difference vectors separately for each target class. This avoids feature entanglement observed in CDCT, where a single VAE is trained on trajectories across all classes. As a result, the extracted directions are more semantically coherent and clinically interpretable.
To further analyze the visual characteristics of the extracted directions, the behavior of the two latent encoders used in the framework is compared. While both encoders support meaningful concept discovery, the CDCT encoder often fails to reconstruct subtle yet clinically relevant features, limiting the fidelity of concept manipulations. In contrast, the LDM encoder preserves fine-grained details and structural consistency, yielding more realistic and diagnostically coherent manipulations. Nonetheless, directions in the LDM space frequently introduce pinkish or violet hues, possibly reflecting biases from the generative prior, which may influence the visual appearance of concepts.
The concept directions discovered by CDLC demonstrate alignment with established dermoscopic features. For instance, directions targeting Melanoma often exhibit asymmetric pigmentation and color variegation, features of malignancy reported in clinical literature [30]. In Basal Cell Carcinoma, the extracted directions introduce translucent reddish or violet hues with dark pigment patches, consistent with features such as ulceration and blue-gray dots [31]. For Dermatofibroma, directions emphasize a central scar-like area of pink hues reported in literature [32], while Vascular Lesions display purplish centers with surrounding pink diffusion, resembling the vascular blush of angiomas [33]. Similarly, directions for Squamous Cell Carcinoma reveal white scaly textures and irregular vascular patterns, characteristic of keratinizing and sun-damaged lesions [34].
In addition to replicating known diagnostic cues, CDLC also uncovers subtler variations such as peripheral pigmentation or central hue changes, that may indicate dataset-specific biases or underexplored clinical markers, requires clinical feedback. In contrast, reliable concept directions could not be identified for the Nevus class. This is likely due to class imbalance, as the dominance of Nevus in the training data results in fewer samples being mapped as counterfactual to it. Consequently, the limited number of latent difference vectors hinders the discovery of consistent semantic directions for this class.
@@ 6 Limitations & Future work
While CDLC demonstrates promising results in extracting semantically meaningful and class-specific concept directions, several limitations remain. The reliability of extracted directions depends on the quality of generated counterfactuals, artifacts or semantic drift in this step may introduce noise into the latent differences. Additionally, spherical clustering assumes that concept directions are well-separated based on angular similarity on a unit hypersphere, which may not adequately capture complex latent geometries or overlapping semantic structures.
Future work could explore alternative encoders, such as Diffusion Autoencoders [35], to enhance reconstruction fidelity and semantic expressiveness. Extending CDLC to other modalities, including multimodal or temporal data, may further broaden its applicability. Finally, incorporating human-in-the-loop evaluation or clinical feedback would support validation and real-world integration of discovered concepts.
@@ 7 Conclusion
This work introduced CDLC, a framework for discovering global, class-specific concept directions by clustering latent differences between factual and counterfactual image pairs. Experiments on a skin lesion classification task demonstrated that the extracted directions influence model predictions and align with known dermoscopic features, supporting their interpretability. However, some concept directions, especially obtained from the LDM encoder, consistently exhibit pinkish or purplish hues, irrespective of the target class. This visual redundancy may indicate an inherent bias in the generative prior, suggesting a need for further analysis to disentangling meaningful concepts from model or dataset-specific artifacts. The ability to uncover diverse, clinically relevant concepts positions CDLC as a valuable tool for advancing concept-based explainability in high-stakes domains.
@@ 10 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used chatgpt in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
11 Data availability
This study uses publicly available dermoscopic images from the ISIC challenges (2016–2020). The exact dataset splits used for training, validation, and testing will be made available upon request.
@@ 11 Data availability
Extraction of Concept Directions Using the CDCT Encoder

### v2

@@ 4.2 Generation of Counterfactuals
The same classifier-guided counterfactual generation approach introduced in the CDCT framework [varshney2025generating] is adopted, with a simplified setting. Rather than capturing intermediate reconstructions across the trajectory, only the final counterfactual image is retained. Counterfactuals are generated using a Latent Diffusion Model based on the Stable Diffusion (SD) 2.1 architecture [^2] Model available at: https://huggingface.co/stabilityai/stable-diffusion-2-1., which is fine-tuned on the consolidated ISIC training dataset to align the generative manifold with the dermoscopic image domain. For counterfactual generation, the same hyperparameters (guidance scale, diffusion steps) are adopted as reported in CDCT, to ensure consistency and comparability (see Appendix A for hyperparameters). For each training sample, counterfactuals are synthesized for all target classes other than the class predicted by the classifier.
@@ 4.3 Extraction of Concept Directions
Semantic concept directions induced by classifier-guided counterfactuals are explored in two distinct latent spaces: (1) the pretrained encoder of the Stable Diffusion 2.1 model, referred to as the LDM encoder, and (2) a Variational Autoencoder trained on counterfactual trajectories, referred to as the CDCT encoder (see Step 2 of the CDCT framework [varshney2025generating] for further details). Experimental setup, results, and analysis based on the CDCT encoder are provided in the Supplementary Material (Appendix E, F).
@@
The unit vectors are collected for all training samples that are not predicted as a given target class. Spherical K-Means clustering [hornik2012spherical] is then applied to group these vectors that reflect similar semantic transformations. The number of clusters $K$ is selected based on the highest silhouette score (see Appendix A for per-class $K$ values). A detailed ablation study analyzing the effect of the number of clusters $K$ is provided in the Supplementary Material (Appendix B). Each resulting cluster represents a set of consistent latent transformations, and the average unit direction within a cluster is interpreted as a representative concept direction for the target class.
@@ 4.4 Interpretation and Evaluation of Concept Directions
To assess the interpretability and effectiveness of the extracted concept directions, each identified concept direction is applied to unseen test images in the latent space. A detailed ablation study analyzing the effect of the scaling parameter $\alpha$ on concept traversal is provided in the Supplementary Material (Appendix C). Figure 2 illustrates concept directions identified by CDLC using the LDM encoder. For Melanoma, the extracted direction (Figure 2a) increases central pigmentation and introduces a purplish hue in the core lesion. In Basal Cell Carcinoma, detected concept direction (Figure 2b) produces a reddish-purple hue with blue-gray dots. For Benign Keratosis, a concept direction (Figure 2c) induces purplish blotches, irregular pigmentation, and blurred borders. In the case of Dermatofibroma, the direction (Figure 2d) introduces a violet-toned center with subtle border irregularities and increased asymmetric pigmentation. For Vascular Lesions, the direction (Figure 2e) enhances central coloration with purplish tones and peripheral pink diffusion. Finally, Squamous Cell Carcinoma directions (Figure 2f) reveal white central structures accompanied by irregular linear streaks suggestive of vascular features.
Table 1 reports the quantitative results for the detected concept directions. Success Rate (SR) measures the proportion of test samples where traversal along the direction increases the probability of the target class, relative to the reconstructed image. LPIPS [zhang2018unreasonable] and FID [heusel2017gans] assess perceptual and distributional realism, while TCAV quantifies concept alignment with the model’s decision boundary, computed from the final conv3 outputs of the third bottleneck block in layer 4 of ResNet50. A detailed analysis of TCAV computed across different layers of ResNet50 for these concept directions can be found in the Supplementary Material (Appendix D). Most concepts achieve high SR (70–85%), confirming their strong influence on classifier decisions. Concepts such as Central Pinkish Veil and Central Pinkish Veil with Asymmetric Intensified Pigmentation attain both high SR and near-perfect TCAV, indicating consistent use by the model. In contrast, White Structures with Irregular Vessels show lower SR and TCAV, suggesting weaker alignment. Overall, SR, LPIPS, FID, and TCAV demonstrate that the discovered concepts are semantically meaningful and predictive of model behavior.
Table 1: Quantitative results for the detected concept directions, reported in terms of Success Rate (SR), LPIPS, FID, and TCAV. Together, these metrics demonstrate that the discovered concepts are both semantically meaningful and predictive of model behavior.
Concept SR (%) $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$ TCAV $\uparrow$
Purplish Core Pigmentation 70.5 0.12 30.5 0.97
Reddish Core with Blue-Gray Dots 83.6 0.17 43.1 0.82
Blotchy Pigmentation with Irregular Texture 81.1 0.20 47.0 0.92
Central Pinkish Veil with Asymmetric Intensified Pigmentation 85.4 0.15 47.1 1.00
Central Purplish Veil 84.6 0.15 53.5 1.00
White Structures with Irregular Vessels 70.5 0.18 51.2 0.64
4.5 Computational Efficiency
[…]
Method Hardware Time Storage
CDCT (trajectory) GPU 1d 16h 25m 314 MB
CDLC (embed. diff.) GPU 1d 18h 35m 69 MB
CDCT (dim. search) GPU $\sim$80 m
CDLC (clustering) CPU $\sim$15 m
@@ 5 Discussion
[…]
To further analyze the visual characteristics of the extracted directions, the behavior of the two latent encoders used in the framework is compared. While both encoders support meaningful concept discovery, the CDCT encoder often fails to reconstruct subtle yet clinically relevant features, limiting the fidelity of concept manipulations. In contrast, the LDM encoder preserves fine-grained details and structural consistency, yielding more realistic and diagnostically coherent manipulations. Nonetheless, directions in the LDM space frequently introduce pinkish or violet hues, which may influence the visual appearance of concepts. Possible explanations include the fact that CDLC learns directions in latent difference space and does not explicitly optimize for hue variations; such hue shifts may therefore arise from generative priors in the latent-diffusion model, from color correlations learned by the target classifier, or from relevant color biomarkers. Therefore, color changes alone are not interpreted as concepts; instead, the qualitative analysis emphasizes structural and textural cues that consistently co-occur with the discovered directions.
The concept directions discovered by CDLC demonstrate alignment with established dermoscopic features. For instance, directions targeting Melanoma often exhibit asymmetric pigmentation and color variegation, features of malignancy reported in clinical literature [argenziano2003dermoscopy]. In Basal Cell Carcinoma, the extracted directions introduce translucent reddish or violet hues with dark pigment patches, consistent with features such as ulceration and blue-gray dots [altamura2010dermatoscopy]. For Dermatofibroma, directions emphasize a central scar-like area of pink hues reported in literature [agero2006conventional], while Vascular Lesions display purplish centers with surrounding pink diffusion, resembling the vascular blush of angiomas [piccolo2018dermatoscopy]. Similarly, directions for Squamous Cell Carcinoma reveal white scaly textures and irregular vascular patterns, characteristic of keratinizing and sun-damaged lesions [rosendahl2012dermoscopy].
In addition to replicating known diagnostic cues, CDLC also uncovers subtler variations such as peripheral pigmentation or central hue changes, which may indicate dataset-specific biases or underexplored clinical markers, and requires clinical feedback. In contrast, reliable concept directions could not be identified for the Nevus class. This is likely due to class imbalance, as the dominance of Nevus in the training data results in fewer samples being mapped as counterfactual to it. Consequently, the limited number of latent difference vectors hinders the discovery of consistent semantic directions for this class.
@@ 6 Limitations & Future work
While CDLC demonstrates promising results in extracting semantically meaningful and class-specific concept directions, several limitations remain. The reliability of extracted directions depends on the quality of counterfactuals, as artifacts or semantic drift at this stage can introduce noise into the latent differences. Some concept directions may exhibit hue shifts (e.g., pink/violet), potentially reflecting generative priors, color correlations in the classifier, or potential color biomarkers. Disambiguation of these causes would require targeted controls (e.g., color-invariant rendering, grayscale analyses) and clinical validation, which we consider beyond the scope of this work and leave for future work. Additionally, spherical clustering assumes that concept directions are well-separated based on angular similarity on a unit hypersphere, which may not adequately capture complex latent geometries or overlapping semantic structures. Exploring more flexible clustering approaches (e.g., kernelized or manifold-based methods) remains an avenue for future work.
Future work could explore alternative encoders, such as Diffusion Autoencoders [preechakul2022diffusion], to enhance reconstruction fidelity and semantic expressiveness. Extending CDLC to other modalities, validating its transferability across domains, and testing it with alternative architectures may further broaden its applicability. Finally, incorporating human-in-the-loop evaluation or clinical feedback would be essential to establish the medical validity of the discovered concepts and support their real-world integration. Such validation requires substantial time and collaboration with domain specialists and is left for future work.
@@ 7 Conclusion
This work introduced CDLC, a framework for discovering global, class-specific concept directions by clustering latent differences between factual and counterfactual image pairs. Experiments on a skin lesion classification task demonstrated that the extracted directions influence model predictions and align with known dermoscopic features, supporting their interpretability. However, some concept directions, especially obtained from the LDM encoder, consistently exhibit pinkish or purplish hues, irrespective of the target class. This visual redundancy may indicate an inherent bias in the generative prior, suggesting a need for further analysis to disentangle meaningful concepts from model or dataset-specific artifacts. The ability to uncover diverse, clinically relevant concepts positions CDLC as a valuable tool for advancing concept-based explainability in high-stakes domains.
@@ 10 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used GPT-4o in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
@@
Appendix A Hyperparameters
For counterfactual generation, we followed the CDCT [varshney2025generating] framework and adopted the same hyperparameter settings: guidance scale = 4 and number of diffusion steps = 10. The number of clusters $K$ was determined using the highest silhouette score. The selected $K$ values for each class are reported in Table 3.
Table 3: Number of clusters $K$ selected for each skin lesion class using highest the silhouette score.
Class Clusters $K$
Melanoma 8
Nevus 5
Basal Cell Carcinoma 6
Benign Keratosis 7
Actinic Keratosis 6
Dermatofibroma 7
Vascular Lesions 6
Squamous Cell Carcinoma 5
Appendix B Ablation Study on the Number of Clusters $K$
For each lesion class $y$, the number of clusters $K$ is determined using the silhouette score. To validate this choice, an ablation study is performed by varying $K\in\{4,5,6,7,8,9\}$ and measuring the downstream effects and the diversity of the learned concept directions.
Let $X$ denote the test set. For class $y$ and image $x\in X$, the $k$-th concept direction (cluster center) is denoted by $c_{y,k}$, and the model probability for class $y$ by $f_{y}(\cdot)$. Applying a concept with scaling $\alpha$ produces a transformed input
$x_{\alpha,k}\;=\;T(x;\,\alpha,\,c_{y,k}),$
and the corresponding per–concept effect is defined as
$\Delta_{y}(x,k)\;=\;f_{y}(x_{\alpha,k})\;-\;f_{y}(x).$
Coverage. Coverage measures the fraction of images for which at least one concept produces a positive effect above a small threshold $\delta>0$ (e.g., $\delta=0.05$):
$\mathrm{Coverage}_{y}(K)\;=\;\frac{1}{|X|}\sum_{x\in X}\mathbbm{1}\!\left[\max_{1\leq k\leq K}\Delta_{y}(x,k)\,\geq\,\delta\right].$
This metric captures the availability of at least one relevant concept for a given image.
Best-of-$K$ influence. The Best-of-$K$ influence measures the gain per image when the most relevant concept is applied:
$\mathrm{BestInfluence}_{y}(K)\;=\;\frac{1}{|X|}\sum_{x\in X}\max_{1\leq k\leq K}\Delta_{y}(x,k).$
This metric serves as an upper bound on actionable improvement, without requiring any single concept to be effective for all images.
Robust mean influence (top-$q$ mean). A robust average is computed to reduce the impact of weak concepts. For $q\in(0,1]$ (with $q=0.3$ in our experiments) and $m=\lceil qK\rceil$, the top-$q$ mean is
$\mathrm{TopQMean}_{y}(K)\;=\;\frac{1}{|X|}\sum_{x\in X}\frac{1}{m}\sum_{k\in\mathrm{top}_{m}\{\Delta_{y}(x,1{:}K)\}}\Delta_{y}(x,k).$
This measure captures the contribution of several useful concepts per image while limiting the influence of outliers.
Redundancy index (diversity). A redundancy index quantifies diversity among concept directions. With unit–norm directions $\hat{c}_{y,k}$ for $k=1,\dots,K$, redundancy is defined as the mean pairwise cosine similarity between all distinct pairs of directions:
$\mathrm{Redundancy}_{y}(K)\;=\;\frac{2}{K(K-1)}\sum_{1\leq i<j\leq K}\big\langle\hat{c}_{y,i},\,\hat{c}_{y,j}\big\rangle.$
Lower values indicate more diverse (less duplicated) concepts, while orthogonal sets yield values near zero.
Table 4 reports these metrics for $K\in\{4,5,6,7,8,9\}$ for each class $y$. Across classes, Coverage, Best-of-$K$, and the top-$q$ mean vary only slightly with $K$, indicating that downstream performance is largely stable with respect to the number of clusters. Best-of-$K$ tends to increase as $K$ grows, while Coverage and the top-$q$ mean show only small, class-specific gains. Redundancy exhibits modest class-specific fluctuations and is often lowest near the silhouette-selected $K$. Although the exact concept directions differ when $K$ changes, their aggregate usefulness and diversity remain stable. This study confirms that the silhouette-based choices reported in the main text (Mel=8, BCC=6, BKL=7, DF=7, VL=6, SCC=5) represent robust operating points: they achieve near-maximal performance while avoiding unnecessary redundancy, and the results are not sensitive to the precise choice of $K$. Moreover, this ablation framework provides a general strategy for selecting an appropriate number of clusters. When the method is transferred to other domains, the same analysis can guide the choice of $K$ in a robust and data-driven manner.
Table 4: Ablation study on the number of clusters $K$ for each lesion class $y$. Reported metrics: Redundancy Index (lower is better), Coverage, Best-of-$K$, and Robust Mean Influence (top-$q$ mean, q=0.3q{=}0.3). Boldface in the $K$ column marks the silhouette-selected $K$; boldface within metric columns marks per-class best values.
Target Class Number of Clusters $K$ Redundancy Index Coverage Best-of-K Influence Robust Mean Influence (top-q mean)
Melanoma 4 0.69 0.72 16.7 13.7
5 0.72 0.75 19.1 15.8
6 0.73 0.76 20.7 17.3
7 0.74 0.77 21.1 15.6
8 0.68 0.78 21.8 16.2
9 0.69 0.79 22.3 17.3
Basal Cell
Carcinoma
4 0.47 0.71 20.9 17.2
5 0.42 0.70 20.0 17.2
6 0.39 0.69 20.3 17.4
7 0.37 0.70 21.1 15.6
8 0.41 0.74 23.2 17.6
9 0.41 0.75 24.2 19.0
Benign
Keratosis
4 0.46 0.84 39.0 31.0
5 0.48 0.85 40.4 33.4
6 0.42 0.85 40.2 33.7
7 0.43 0.85 41.0 29.7
8 0.41 0.86 42.9 32.2
9 0.43 0.87 44.8 34.3
Dermatofibroma 4 0.62 0.53 10.0 5.34
5 0.70 0.55 11.2 6.31
6 0.72 0.57 11.7 7.75
7 0.57 0.57 11.7 4.62
8 0.59 0.60 12.2 5.44
9 0.61 0.60 13.2 5.60
Vascular
Lesions
4 0.68 0.62 19.1 13.4
5 0.60 0.62 19.1 13.4
6 0.62 0.64 20.7 15.2
7 0.57 0.64 20.8 11.9
8 0.65 0.68 24.0 16.1
9 0.62 0.70 24.7 16.8
Squamous Cell
Carcinoma
4 0.43 0.40 3.55 2.09
5 0.39 0.41 3.96 2.50
6 0.46 0.47 4.98 3.22
7 0.42 0.48 5.58 2.62
8 0.43 0.49 5.88 2.94
9 0.37 0.45 5.17 2.70
Appendix C Ablation Study on the scaling parameter $\alpha$
To evaluate the sensitivity of the scaling parameter $\alpha$ for the detected concepts, experiments are conducted across representative concepts. Table 5 reports the effect of varying $\alpha$ on three metrics: Success Rate (SR), which indicates the proportion of test samples for which traversal along the corresponding direction increases the predicted probability of the target class relative to the reconstructed image, LPIPS [zhang2018unreasonable] measures perceptual realism of generated images (lower is better), and FID [heusel2017gans] assesses overall image distributional fidelity (lower is better).
The results reveal a trade-off: smaller $\alpha$ values yield more realistic images (lower LPIPS/FID) but weaker concept manifestation (lower SR), whereas larger $\alpha$ values improve SR but may introduce artifacts and reduce realism. The optimal $\alpha$ for each concept is highlighted in bold and reflects a balanced trade-off between edit success and perceptual fidelity.
Notably, the best-performing $\alpha$ varies across concepts (e.g., $40$ for Purplish Core Pigmentation, and $70$ for Reddish Core with Blue-Gray Dots), confirming that a single global value of $\alpha$ is suboptimal. For most concepts, moderate $\alpha$ values provide a stable regime where SR increases steadily with only minor degradation in perceptual metrics, whereas higher values often exaggerate the edits, leading to artifacts despite marginal SR gains.
Table 5: Ablation study on the effect of $\alpha$ for selected concept directions. The best values for each metric are highlighted in bold. The chosen $\alpha$ (bold entry in the first column) represents the optimal balance between concept success rate and perceptual fidelity. Lower LPIPS and FID values indicate higher perceptual quality and image realism.
Concept $\boldsymbol{\alpha}$ SR (%) $\uparrow$ LPIPS $\downarrow$ FID $\downarrow$
Purplish Core Pigmentation 40 70.54 0.123 30.51
45 69.86 0.130 35.56
50 68.34 0.136 40.42
55 66.23 0.143 45.08
60 63.45 0.150 49.77
Reddish Core with Blue-Gray Dots 50 74.68 0.138 22.11
55 77.94 0.145 26.58
60 80.31 0.153 31.73
65 81.84 0.160 37.10
70 83.61 0.168 43.05
Blotchy Pigmentation with Irregular Texture 70 72.61 0.165 29.10
75 75.36 0.173 33.12
80 78.11 0.181 37.65
85 79.50 0.189 42.13
90 81.09 0.198 47.01
Central Pinkish Veil with Asymmetric Intensified Pigmentation 40 80.28 0.124 32.04
45 82.35 0.131 35.86
50 83.88 0.137 39.54
55 84.96 0.143 43.29
60 85.44 0.149 47.13
Central Purplish Veil 40 76.55 0.125 34.61
45 79.13 0.132 39.63
50 81.26 0.138 44.31
55 82.69 0.145 48.90
60 84.59 0.152 53.48
White Structures with Irregular Vessels 60 51.54 0.148 31.35
65 56.83 0.155 36.14
70 62.50 0.163 40.98
75 66.30 0.170 45.98
80 70.47 0.177 51.21
Appendix D Evaluation of concept directions: TCAV score
Table 6: TCAV scores for representative concepts across different layers of the ResNet50 architecture. Higher scores indicate stronger alignment between the concept direction and the model’s decision boundary. Bold entries denote the layer with the highest score for each concept. Here, layer4-B0 (conv3), layer4-B1 (conv3), and layer4-B2 (conv3) refer to the final conv3 outputs of the first, second, and third bottleneck blocks in the layer 4, respectively.
Concept layer1 layer2 layer3 layer4-B0 (conv3) layer4-B1 (conv3) layer4-B2 (conv3)
Purplish Core Pigmentation $0.46\pm 0.04$ $0.35\pm 0.06$ $0.31\pm 0.13$ $0.39\pm 0.10$ $0.77\pm 0.02$ $\mathbf{0.97\pm 0.02}$
Reddish Core with Blue-Gray Dots $0.73\pm 0.01$ $\mathbf{0.93\pm 0.01}$ $0.89\pm 0.03$ $0.77\pm 0.07$ $0.82\pm 0.07$ $0.82\pm 0.13$
Blotchy Pigmentation with Irregular Texture $0.64\pm 0.02$ $0.76\pm 0.01$ $0.73\pm 0.10$ $0.72\pm 0.08$ $0.86\pm 0.07$ $\mathbf{0.92\pm 0.02}$
Central Pinkish Veil with Asymmetric Intensified Pigmentation $0.70\pm 0.01$ $0.96\pm 0.01$ $0.88\pm 0.01$ $0.95\pm 0.02$ $\mathbf{1.00\pm 0.00}$ $\mathbf{1.00\pm 0.00}$
Central Purplish Veil $2\pm 0.02 0.77±0$$7\pm 0.02$ $0.96\pm 0.02$ $0.98\pm 0.01$ $0.98\pm 0.00$ $\mathbf{1.00\pm 0.00}$
White Structures with Irregular Vessels $\mathbf{0.76\pm 0.01}$ $0.66\pm 0.00$ $5\pm 0.00 0.66±0$$6\pm 0.06$ $0.74\pm 0.10$ $0.64\pm 0.18$
Table 6 reports the TCAV [kim2018interpretability] scores computed across different layers of ResNet50 [he2016deep] (target model) architecture for representative concepts. Higher TCAV values indicate a stronger alignment between the concept direction and the model’s decision boundary.
Several insights emerge from these results. First, concept importance varies substantially across layers, highlighting that semantic information is not uniformly distributed throughout the network. For example, Purplish Core Pigmentation exhibits a clear progression from low scores in early layers (0.46 in layer1) to a very strong alignment in deeper layers (0.97 at the final conv3 of the second bottleneck in layer4), suggesting that this concept is primarily captured at higher semantic levels. Similarly, Blotchy Pigmentation with Irregular Texture and Central Purplish Veil both achieve their peak TCAV scores in deeper blocks, indicating that these nuanced features emerge late in the representation hierarchy.
In contrast, Reddish Core with Blue-Gray Dots reaches its maximum importance at an intermediate layer (0.93 in layer2), suggesting that certain concepts are best represented at mid-level feature abstraction. Interestingly, White Structures with Irregular Vessels shows the opposite trend, with its highest score at the shallowest layer (0.76 in layer1), implying that this concept is captured through low-level visual cues such as edge or color contrasts.
These findings demonstrate that different clinical concepts are encoded at different depths of the network: some emerge only in higher-level semantic layers, while others are already salient at earlier stages. This underlines the importance of evaluating multiple layers when applying concept-based interpretability methods.
Appendix E Extraction of Concept Directions Using the CDCT Encoder